OpenStudy (anonymous):

What is 3 · 2 1/3 ? A. 7 B. 6 1/3 C. 6 D. 5 1/3

OpenStudy (anonymous):

@chosenmatt

OpenStudy (igreen):

Change 2 1/3 into an improper fraction. Multiply the denominator to the whole number, then add it to the numerator. \(3 \times 2 + 1 = ?\) @THEUSER822

OpenStudy (chosenmatt):

one sec

OpenStudy (anonymous):

7

OpenStudy (igreen):

Yep!

OpenStudy (igreen):

That gives us: \(\dfrac{3}{1} \times \dfrac{7}{3}\) \(\dfrac{3 \times 7}{1 \times 3}\) Can you multiply those? @THEUSER822

OpenStudy (anonymous):

yes @iGreen

OpenStudy (anonymous):

\[\frac{ 21 }{ 3 }\]

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

And can you divide 21 / 3? @THEUSER822

OpenStudy (anonymous):

7